OpenStudy (anonymous):

what is "initial velocity" and how do we find it? Like whats the difference between that and "average velocity"

OpenStudy (anonymous):

Initial velocity means the velocity with which the body started. Initial velocity when the object starts from rest is 0

OpenStudy (ash2326):

Initial velocity is the velocity of a moving body at the starting. if we are given final velocity, acceleration and the time for which the body is accelerated we can find the initial velocity. v-u=at u=v-at

OpenStudy (anonymous):

Make u the subject of the formula in v=u+at which is a physics eqn. v=final velocity u=initial velocity a=acceleration t=time ie if you have the variables

OpenStudy (ash2326):

while Average velocity is the displacement/time

OpenStudy (anonymous):

ok so if the question is asking what Find the initial velocity for an object that moves along a straight line so that at time "t" measured in seconds, its distance from a fixed point O is given by x= 8 -2t -t^2

OpenStudy (ash2326):

differentiate this to find velocity x=8-2t-t^2 dx/dt=-2-2t v=-2-2t

OpenStudy (ash2326):

put t=0 to get initial velocity v=-2 initial velocity

OpenStudy (anonymous):

ohhhh i seeee arigato

OpenStudy (ash2326):

welcome